# OutLines: Modeling Spectral Lines from Winds, Bubbles, and Outflows

### v1

@@ Abstract
Common methods for studying the kinematics and geometry of outflowing gas rely on modeling emission and absorption lines in integrated spectra using methods that are not physically motivated, including empirical quantiles or fitting multiple Gaussian or Voigt profiles. Such methods are not always consistent with the interpretation of these features and, as a result, miss key underlying physics and can even lead to inaccurate interpretations of observations. To address this problem, we present the publicly available python code OutLines, which provides astrophysical models of spectral emission and absorption line profiles produced by outflows in a variety of environments. The OutLines code accounts for differences in parameterization of the velocity field and density profile while allowing for different outflow geometries, making OutLines versatile and useful for a wide variety of astrophysical phenomena. We demonstrate the wide applicability of OutLines by using the code to model line profiles in an H ii region knot, super star clusters, a starburst galaxy, and an AGN. In each of these contexts, we illustrate how OutLines can illuminate key underlying physics in ways that improve our scientific understanding and address important open questions in astronomy, including the key mechanisms in the baryon cycle, the evolution of H ii regions and galaxies, and even Lyman continuum escape. OutLines will be a critical resource as massively multiplexed spectroscopic surveys like WEAVE-LOFAR and 4MOST/WAVES come online, providing the means to probe feedback kinematics with deeper, higher resolution spectroscopy for unprecedented large samples of galaxies.
@@
pubyear: 2025pagerange: OutLines: Modeling Spectral Lines from Winds, Bubbles, and Outflows28
@@ 1 Introduction
Outflows, winds, and bubbles are ubiquitous astrophysical phenomena occurring on a dynamic range of spatial scales and in a variety of contexts, from young stellar objects (YSOs) to nebulae, active galactic nuclei (AGN), galaxies, and even galaxy clusters. This feedback has the capacity to shape many key astrophysical systems, from YSO winds affecting the formation of exoplanets (e.g., Suzuki2009; Manara2023) to massive stars redistributing the leftover material from their birth clouds (e.g., Dopita2005; Menon2024) to AGN and starbursts driving the baryon cycle of galaxies (e.g., 2005ApJ...618..569M; Tumlinson2017). Many physical mechanisms can accelerate, or drive, winds and bubbles, from cosmic rays (e.g., Ipavich1975) to hot expanding gas (e.g., 1985Natur.317...44C) to radiation pressure on dust (e.g., Poynting1904; LyndenBell1971) and ions (so-called line driving, e.g., Lucy1970), each with direct implications for how feedback shapes the surrounding environment (see, e.g., 2020A&ARv..28....2V; Thompson2024, and references therein). A common method of detecting and studying feedback is through broad, often asymmetric spectroscopic features (e.g., Castor1970; 1975ApJ...195..157C; 1974ApJ...193..651N; 1980A&A....81..172P; 1981MNRAS.195..787P; 1981ApJ...247..403H; Chu1986), perhaps most famously with P Cygni wind lines in stellar atmospheres (e.g., 1986A&A...164...86P; 1987ApJ...314..726L; 2001A&A...369..574V). Physically motivated models for these line profiles is thus of critical importance to understanding how energetic phenomena shape their environment.
@@
While the physics and formalism for spectroscopic features arising from winds, bubbles, and outflows is not new (e.g., Beals1931; 1960mes..book.....S; Castor1970; 1975ApJ...195..157C; 1986A&A...164...86P; 1987ApJ...314..726L), their application is relatively sparse outside the phenomena of stellar atmospheres (e.g., 2001A&A...369..574V; 2021MNRAS.504.2051V; Sander2017; Puls2020). Instead, fitting multiple Gaussian or Voigt profiles to broad, asymmetric lines has become the standard (e.g., 1980A&A....81..172P; 1981MNRAS.195..787P) for a wide variety of objects, including H ii regions (e.g., Bresolin2020; Yarovova2025), AGN (e.g., Liu2020), and galactic winds (e.g., Hogarth2020; Peng2023; Peng2025), despite early evidence that physics-based profiles are not only more appropriate but also demonstrably different in shape while still providing good descriptions of observations (e.g., Zheng1990; TenorioTagle1996). Some studies have skirted the issue by adopting a more empirical approach using velocity quantiles (e.g., 1981ApJ...247..403H; 1985MNRAS.213....1W; 1991ApJS...75..383V) or by fitting the line profile wings with a log-linear (e.g., 2021ApJ...920L..46K; Komarova2025) or exponential (e.g., Rusakov2025) functions to characterize the steepness of the non-Gaussian flux distribution. The benefit of such methods is that they require no physical model while still providing insight into kinematics. However, bridging the gap between these approaches and the underlying physics is not always straightforward.
@@
Some effort towards a physically motivated approach has emerged in recent years, including detailed treatment of radiation pressure and conservation effects (e.g., 2017MNRAS.471.4061K), turbulent mixing layers (e.g., 2009A&A...500..817B), and relativistic accretion disk winds (e.g., DISKWIND1; DISKWIND2; XRADE) along with more generalized models (e.g., 2010ApJ...717..289S; 2011ApJ...734...24P; Martin2013; 2015ApJ...801...43S; 2016MNRAS.463..541C; Luminari2018; Flury2023). Despite these exciting advancements, wide-spread adoption of these methods has stalled. A key to furthering the use of physically-motivated line profiles is versatile, well-documented, and user-friendly software tools which make such models readily accessible to the community to generate spectral line profiles linking observations to outflow physics.
@@
In this paper, we present OutLines, a python-based code for generating model emission, absorption, and P Cygni line profiles arising from astrophysical winds, bubbles, and outflows. The input geometry and wind physics can be readily customized, and functionality for modeling is provided. We outline the formalism and physics for radiative transfer, velocity fields, gas density profiles, and geometry in §2. Then, in §3 we present the implementation of these components in the OutLines code111available at github.com/sflury/OutLines and via pip install.. We demonstrate the versatile applicability and scientific usefulness of OutLines in §4 by modeling broad spectral lines in an H ii region knot, a super star cluster (SSC), a starbursting galaxy, and an active galactic nucleus (AGN), with profiles generated by OutLines. Finally, we summarize OutLines and its potential for high scientific impact in §5, providing worked examples in Appendix D and E.
@@ 2 Physical Model
Within the framework of the Sobolev approximation (1944SvA.....21..143S; 1957SvA.....1..678S; 1960mes..book.....S), small scale Doppler motions of gas within clouds can be ignored in the context of the large-scale motion of the clouds themselves. The Sobolev approach has a long-standing precedent in the study of winds (e.g., Castor1970; 1974ApJ...193..651N; 1975ApJ...202..465M; 1987ApJ...314..726L), with successful application to galactic winds and outflows over the past 15 years (e.g., 2010ApJ...717..289S; 2011ApJ...734...24P; Martin2013; 2015ApJ...801...43S; 2016MNRAS.463..541C; Flury2023). Under this assumption, the key characteristics of winds, outflows, and bubbles which dictate the line profile are the velocity field, density profile, and geometry of the clouds (e.g., Zheng1990). Below, we elaborate how the observed projected velocity $u$ in a spectral line, given by
$u=\frac{\lambda-\lambda_{0}}{\lambda_{0}}\frac{c}{v_{\infty}}$ (1)
for a line with rest wavelength $\lambda_{0}$ and normalized to the wind terminal velocity $v_{\infty}$, relates to the intrinsic wind velocity $w$ given by some radially-symmetric velocity field
@@ 2.1.1 Line Absorption
Under the Sobolev approximation, instead one must evaluate the relative column density at each velocity $v$ in the wind or bubble that is projected onto the line of sight velocity $u$. The optical depth profile for the outflow component is given by the Sobolev approximation such that
$\phi_{\lambda}=K^{-1}\int n[x]\frac{{\rm d}x[w]}{{\rm d}w}{\rm d}\Omega$ (5)
over some differential solid angle ${\rm d}\Omega$ given by the geometry (see Rybicki1983, their Equation 2.14) and where $K$ is a normalization constant ensuring $\int_{\lambda}\phi_{\lambda}{\rm d}\lambda=1$ (Rybicki1979, their Equation 1.65). The differential term $\frac{{\rm d}x}{{\rm d}w}$ arises from integrating over all column densities associated with the observed velocity $u$ (Castor1970, their Equation 6, 2016MNRAS.463..541C, their Equation 4, assuming no stimulated emission). The corresponding integral limits are given by constraints from the geometry (§2.4).
@@ 2.1.2 Nebular Line Emission
for any given wavelength. “Static” gas undergoes Doppler broadening from thermal, orbital, and turbulent motion which produces the traditional Gaussian form of $\phi_{\lambda}$. In the case of an outflow, we must first consider how the line emission depends on density and velocity. The line emission coefficient $j_{\lambda}$ is typically defined as proportional to the density of the populated upper state producing the transition (Rybicki1979; source:osterbrock2006). Since nebular lines emerge from upper electronic states populated by recombination or collisional excitation, $j\propto n[x(w)]^{2}$ (e.g., source:osterbrock2006; Peimbert_2017). Thus, we can simply express the line profile as
$\phi_{\lambda}=K^{-1}\int n\left[x[w]\right]^{2}{\rm d}\Omega$ (10)
where $K$ again is a normalization constant ensuring $\int_{\lambda}\phi_{\lambda}{\rm d}\lambda=1$ (see Flury2023). As with absorption, limits on the integral are given by the geometry in §2.4.
@@ 2.1.3 Resonant and Fluorescent Line Emission
Unlike nebular line emission, which depends on collisional and recombination processes, the upper state population of resonant and fluorescent lines depends primarily on the absorption of photons which are subsequently re-emitted in a random direction. Therefore, the line emission coefficient $j_{\lambda}$ is proportional to nndx/x/dww, meaning that the line profile is determined using Equation 5 but with integral limits for line emission given by integrating over the entire volume (see §2.4). However, photons emitted in this fashion can be re-absorbed and subsequently re-emitted in an arbitrary direction, effectively a scattering process, which will alter the line profile. Resonant and fluorescent lines re-emit light which has previously been absorbed isotropically, meaning that the line flux is determined by the source function $S_{\lambda}=1-\exp[-\tau_{\lambda}]$ (see, e.g., Castor1970, their §6, Rybicki1983, their §3) such that
@@
where $\tau_{\lambda}$ is given by line absorption in Equations 4 and 5 with integral limits consistent with line emission in Equation 10, $\epsilon$ is the filling factor to account for clumpiness because individual clouds do not sample every possible sightline even when they are isotropically distributed, and $f_{\rm esc,\lambda}$ is the fraction of photons which escape radiative trapping at a given wavelength $\lambda$. The escape fraction of photons depends on the fraction $p_{i}=A_{i}/\sum A$ of photons which are produced by a transition $i$ given by the transition rate $A_{i}$ divided by the total radiative depopulation rate $\sum A$, where each $i$ corresponds to a resonant ($p_{res}$) or fluorescent ($p_{fluor}$) channel. Following formalism from Castor1970, the probability that a resonant photon can escape from radiative “trapping” through resonant scattering is given by the ratio of emission to absorption $(1-\exp[-\tau_{\rm res,\lambda}])/\tau_{\rm res,\lambda}$. Trapped photons either scatter through emission and re-absorption in the resonant channel or re-emit in the fluorescent channel. After multiple scatterings, the resonant photons are increasingly trapped relative to the fluorescent photons, meaning that the escape fraction is
@@
for fluorescent photons under the infinite scattering limit given by 2015ApJ...801...43S.
@@ 2.2 Solutions to Wind Equations of Motion
Velocity profiles for winds can take on a variety of shapes arising from different treatments of the wind equations of motion (e.g., CAK theory 1975ApJ...195..157C; 1976ApJS...32..715L; 1977PhDT.........3A; 1977ApJ...213..737B; Castor1979). The simple $\beta$ law approximation
@@
first appears in early treatment of Wolf-Rayet P Cygni lines for $\beta=0.5$ (Equation 23 Castor1970) and arises from a generalization of CAK theory solutions (see, e.g., Equation 47 of 1975ApJ...195..157C) without loss of physical insight (see, e.g., 1977ApJ...213..737B; Castor1979; 1986A&A...164...86P).The CAK $\beta$ law approximation is ubiquitous in the study of stellar winds (e.g., 2001A&A...369..574V; 2021MNRAS.504.2051V; Telford2024) and has even seen application to galactic outflows (e.g., 2016MNRAS.463..541C; Flury2023). Given its widespread adoption and underlying physical motivation, the CAK $\beta$ law is our default and fiducial velocity field.
In addition to the $\beta$ law approximation to CAK theory, OutLines includes two alternate solutions to the wind equations proposed in the literature. The first, from 2010ApJ...717..289S, assumes a power law approximation to the radial acceleration field to obtain
@@
where the normalization terms are implicit in $w$ and $x$ and $r_{min}=1$. While those authors do not provide a corresponding physical model, their acceleration power law has a velocity field qualitatively similar to those of Ipavich1975, which assumed cosmic ray driving, and 1985Natur.317...44C, which assumed constant momentum and energy deposition into a steady-state wind from a fluid in which it is entrained (see also, e.g., Owocki2025), and formally akin to radiative acceleration of expanding clouds (see Equations 16-17 Mathews1974). Note that the acceleration power law field with $\beta=2$ is identical to that of the CAK $\beta$ law when $\beta=0.5$.
The second velocity field alternative, from various sources (e.g., FaucherGiguere2012), constitutes a power law approximation to the radial velocity field such that
@@
where the wind is assumed to accelerate from rest at $R_{0}$ and $A$ scales the field. We set $A=0.5$ to allow clouds to reach 50% terminal velocity in roughly the same range of radii as the CAK theory $\beta$ law and generally overlap with the same region in $w$-$x$ space as both the CAK $\beta$ law and acceleration power law fields. Single-shell models of stellar wind bubbles can give rise to power law velocity fields from the shared time dependence of $r\propto t^{3/5}$ and $v\propto t^{-2/5}$ ($v\propto r^{-2/3}$, e.g., Weaver1977; Dyson1977). Building on this framework, FaucherGiguere2012 arrive at a velocity power law solution assuming continuous production of “shells” by an AGN to produce a wind.
@@
Figure 1 depicts the different velocity fields for a range of $\beta$. From this comparison, as suggested by the formalism, the power law velocity field approaches $v_{\infty}$ more rapidly than the power law acceleration field or the $\beta$-law approximation to CAK theory, reaching $v_{\infty}$ at distances far closer to the launch radius. Moreover, the power law velocity field does not approach $v_{\infty}$ asymptotically as do the other velocity fields, meaning that the stopping condition at $v_{\infty}$ is imposed solely by integral limits rather than the formalism for the velocity field. We note that physical wind models almost always exhibit an acceleration which gradually decreases as the outflow asymptotically approaches some terminal velocity (e.g., 1975ApJ...195..157C; 1977ApJ...213..737B; 1985Natur.317...44C; Thompson2015; 2017MNRAS.471.4061K; 2018MNRAS.476..512I; Quataert2022a), indicating that the underlying wind physics strongly favors a CAK $\beta$ law or acceleration power law field over a velocity power law field. While individual wind bubble models indicate a velocity power law field (e.g., Weaver1977; Dyson1977), in ensembles of shells, the distribution of shell ages will dominate the net velocity field (e.g., TenorioTagle1996) and is likely determined by the physical driving mechanism (e.g., Chu1994), suggesting that the CAK theory or acceleration power law velocity fields provide a more physically appropriate description for wind scenarios. That being said, the velocity power law field prescription can be consistent with turbulent streaming including the classic Kolmogorov cascade, with the case of $\beta=1$ indicative of Couette flow (e.g., 2009A&A...500..817B), although such scenarios may not capture the underlying driving mechanism.
@@
Figure 1: Radial velocity fields included in OutLines and given by Equations 15-17. Acceleration from rest at the base $R_{0}$ of the wind ($x=1$) proceeds until the wind reachs the terminal velocity $v_{\infty}$ ($w[x]=1$). Velocity fields include the 1986A&A...164...86P $\beta$ law approximation to CAK theory (1975ApJ...195..157C) (blue, Equation 15), the 2010ApJ...717..289S power law acceleration (green, Equation 16), and the power law velocity field (purple, Equation 17). Dash length increases with increasing $\beta$.

### v2

@@ Abstract
Studies of kinematics and geometry of outflowing gas rely on modeling features in integrated spectra using empirical quantiles or fitting multiple Gaussian or Voigt profiles. Such methods can miss key underlying physics and even lead to spurious interpretations of observations. To address this problem, we present the public python code OutLines, which predicts spectral emission and absorption line profiles produced by winds, bubbles, and outflows. By design, OutLines is phenomenologically and scale agnostic to bridge the gap between observations and simulations across a broad swath of astrophysics. The OutLines code accounts for differences in velocity field, density profile, and outflow geometries, making OutLines versatile for a wide variety of astrophysical phenomena. We demonstrate the wide applicability of OutLines by using the code to model line profiles in an H ii region knot, super star clusters, a starburst galaxy, and an AGN. In each of these contexts, we illustrate how OutLines can illuminate key underlying physics in ways that improve scientific understanding and address important open questions in astronomy, including the key mechanisms in the baryon cycle, the evolution of H ii regions and galaxies, and even Lyman continuum escape. OutLines will be a critical resource as massively multiplexed spectroscopic surveys with facilities like WEAVE and 4MOST come online, providing the means to probe feedback with deeper, higher resolution spectroscopy for unprecedented large samples of objects.
@@
pubyear: 2026pagerange: OutLines: Modeling Spectral Lines from Winds, Bubbles, and Outflows29
@@ 1 Introduction
Outflows, winds, and bubbles are ubiquitous astrophysical phenomena occurring on a dynamic range of spatial scales and in a variety of contexts, from young stellar objects (YSOs) to nebulae, active galactic nuclei (AGN), galaxies, and even galaxy clusters. This feedback has the capacity to shape many key astrophysical systems, from YSO winds affecting the formation of exoplanets (e.g., Suzuki and Inutsuka, 2009; Manara et al., 2023) to massive stars redistributing the leftover material from their birth clouds (e.g., Dopita et al., 2005; Menon et al., 2024) to AGN and starbursts driving the baryon cycle of galaxies (e.g., Murray et al., 2005; Tumlinson et al., 2017). Many physical mechanisms can accelerate, or drive, winds and bubbles, from cosmic rays (e.g., Ipavich, 1975) to hot expanding gas (e.g., Chevalier and Clegg, 1985) to radiation pressure on dust (e.g., Poynting, 1904; Lynden-Bell and Rees, 1971) and ions (so-called line driving, e.g., Lucy and Solomon, 1970), each with direct implications for how feedback shapes the surrounding environment (see, e.g., Veilleux et al., 2020; Thompson and Heckman, 2024, and references therein). A common method of detecting and studying feedback is through broad, often asymmetric spectroscopic features (e.g., Castor, 1970; Castor et al., 1975a; Noerdlinger and Rybicki, 1974; Pelat and Alloin, 1980; Pelat et al., 1981; Heckman et al., 1981; Chu and Kennicutt, 1986), perhaps most famously with P Cygni wind lines in stellar atmospheres (e.g., Pauldrach et al., 1986; Lamers et al., 1987; Vink et al., 2001). Physically motivated models for these line profiles is thus of critical importance to understanding how energetic phenomena shape their environment.
@@
While the physics and formalism for spectroscopic features arising from winds, bubbles, and outflows is not new (e.g., Beals, 1931; Sobolev, 1960; Castor, 1970; Castor et al., 1975a; Pauldrach et al., 1986; Lamers et al., 1987), their application is relatively sparse outside the phenomena of stellar atmospheres (e.g., Vink et al., 2001; Vink and Sander, 2021; Sander et al., 2017; Puls et al., 2020). Instead, fitting multiple Gaussian or Voigt profiles to broad, asymmetric lines has become the standard (e.g., Pelat and Alloin, 1980; Pelat et al., 1981) for a wide variety of objects, including H ii regions (e.g., Bresolin et al., 2020; Yarovova et al., 2025), AGN (e.g., Liu et al., 2020), and galactic winds (e.g., Hogarth et al., 2020; Peng et al., 2023, 2025), despite early evidence that physics-based profiles are not only more appropriate but also demonstrably different in shape while still providing good descriptions of observations (e.g., Zheng et al., 1990; Tenorio-Tagle et al., 1996). Some studies have skirted the issue by adopting a more empirical approach using velocity quantiles (e.g., Heckman et al., 1981; Whittle, 1985; Veilleux, 1991) or by fitting the line profile wings with a log-linear (e.g., Komarova et al., 2021, 2025) or exponential (e.g., Rusakov et al., 2025) functions to characterize the steepness of the non-Gaussian flux distribution. The benefit of such methods is that they require no physical model while still providing insight into kinematics. However, bridging the gap between these approaches and the underlying physics is not always straightforward.
@@
Some effort towards a physically motivated approach has emerged in recent years, including detailed treatment of radiation pressure and conservation effects (e.g., Krumholz et al., 2017), turbulent mixing layers (e.g., Binette et al., 2009), and relativistic accretion disk winds (e.g., Sim et al., 2008, 2010; Matzeu et al., 2022) along with more generalized models (e.g., Steidel et al., 2010; Prochaska et al., 2011; Martin et al., 2013; Scarlata and Panagia, 2015; Chisholm et al., 2016; Luminari et al., 2018; Flury et al., 2023). Despite these exciting advancements, wide-spread adoption of these methods has stalled. A key to furthering the use of physically-motivated line profiles is versatile, well-documented, and user-friendly software tools which make such models readily accessible to the community to generate spectral line profiles linking observations to outflow physics without forward-modeling simulations.
@@
In this paper, we present OutLines, a python-based code for generating model emission, absorption, and P Cygni line profiles arising from astrophysical winds, bubbles, and outflows on arbitrary spatial scales and independent of the phenomenon. The input geometry and wind physics can be readily customized, and functionality for modeling is provided. We outline the formalism and physics for radiative transfer, velocity fields, gas density profiles, and geometry in §2. Then, in §3 we present the implementation of these components in the OutLines code111available at github.com/sflury/OutLines and via pip install.. We demonstrate the versatile applicability and scientific usefulness of OutLines in §4 by modeling broad spectral lines in an H ii region knot, a super star cluster (SSC), a starbursting galaxy, and an active galactic nucleus (AGN), with profiles generated by OutLines. In the final case, we demonstrate that OutLines can successfully recover wind geometry through comparison with observations. Finally, we summarize OutLines and its potential for high scientific impact in §5, providing details of code setup and installation in Appendix B and worked examples in Appendices C and D.
@@ 2 Physical Model
Within the framework of the Sobolev approximation (Sobolev, 1944, 1957, 1960), small scale Doppler motions of gas within clouds can be ignored in the context of the large-scale motion of the clouds themselves. The Sobolev approach has a long-standing precedent in the study of winds (e.g., Castor, 1970; Noerdlinger and Rybicki, 1974; Mihalas et al., 1975; Lamers et al., 1987), with successful application to galactic winds and outflows over the past 15 years (e.g., Steidel et al., 2010; Prochaska et al., 2011; Martin et al., 2013; Scarlata and Panagia, 2015; Chisholm et al., 2016; Flury et al., 2023). Under this assumption, the key characteristics of winds, outflows, and bubbles which dictate the line profile are the velocity field, density profile, and geometry of the clouds (e.g., Zheng et al., 1990). Below, we elaborate how the observed projected velocity $u$ in a spectral line, given by the longitudinal relativistic Doppler shift (Einstein, 1905, also Sher 1968 their Eqn 4) such that
$u=\frac{c}{v_{\infty}}\frac{(\lambda/\lambda_{0})^{2}-1}{(\lambda/\lambda_{0})^{2}+1}$ (1)
for a line with rest wavelength $\lambda_{0}$ and normalized to the wind terminal velocity $v_{\infty}$ and reducing to the classical Doppler shift $u=c(\lambda/\lambda_{0}-1)/v_{\infty}$ when $v_{\infty}\ll c$, relates to the intrinsic wind velocity $w$ given by some radially-symmetric velocity field
@@ 2.1.1 Line Absorption
Under the Sobolev approximation, instead one must evaluate the relative column density at each velocity $w$ in the wind or bubble that is projected onto the line of sight velocity $u$. The optical depth profile for the outflow component given by the Sobolev approximation is thus
$\phi_{\lambda}=K^{-1}\int n[x]\frac{{\rm d}x[w]}{{\rm d}w}{\rm d}\Omega^{\prime}$ (5)
over some observed differential solid angle ${\rm d}\Omega$ given by the geometry (see Rybicki and Hummer, 1983, their Equation 2.14) and where $K$ is a normalization constant ensuring $\int_{\lambda}\phi_{\lambda}{\rm d}\lambda=1$ (Rybicki and Lightman, 1979, their Equation 1.65). The differential term $\frac{{\rm d}x}{{\rm d}w}$ arises from integrating over all column densities associated with the observed velocity $u$ (Castor 1970, their Equation 6, Chisholm et al. 2016, their Equation 4, assuming no stimulated emission). The corresponding integral limits are given by constraints from the geometry (§2.4). Following the equation of radiative transfer, one must sum up contributions of each velocity slice to the optical depth, not to the emergent flux, or else risk falling victim to Jensen’s inequality ($\exp[-\int\tau]<\int\exp[-\tau]$) by over-estimating the column density and optical depth (e.g., Jennings et al., 2025).
@@ 2.1.2 Nebular Line Emission
for any given wavelength. “Static” gas undergoes Doppler broadening from thermal, orbital, and turbulent motion which produces the traditional Gaussian form of $\phi_{\lambda}$. In the case of an outflow, we must first consider how the line emission depends on density and velocity. The line emission coefficient $j_{\lambda}$ is typically defined as proportional to the density of the populated upper state producing the transition (Rybicki and Lightman, 1979; Osterbrock and Ferland, 2006). Since nebular lines emerge from upper electronic states populated by recombination or collisional excitation, $j\propto n[x(w)]^{2}$ (e.g., Osterbrock and Ferland, 2006; Peimbert et al., 2017). Thus, we can simply express the line profile as
$\phi_{\lambda}=K^{-1}\int n\left[x[w]\right]^{2}{\rm d}\Omega^{\prime}$ (10)
where $K$ again is a normalization constant ensuring $\int_{\lambda}\phi_{\lambda}{\rm d}\lambda=1$ (see Flury et al., 2023). As with absorption, limits on the integral are given by the geometry in §2.4.
@@ 2.1.3 Resonant and Fluorescent Line Emission
Unlike nebular line emission, which depends on collisional and recombination processes, the upper state population of resonant and fluorescent lines depends primarily on the absorption of photons which are subsequently re-emitted in a random direction. Therefore, the line emission coefficient $j_{\lambda}$ is proportional to nndx/x/dww, meaning that the line profile is determined using Equation 5 but with integral limits for line emission given by integrating over the entire volume (see §2.4). However, photons emitted in this fashion can be re-absorbed and subsequently re-emitted in an arbitrary direction, effectively a scattering process, which will alter the line profile. Resonant and fluorescent lines re-emit light which has previously been absorbed isotropically, meaning that the line flux is determined by the source function $S_{\lambda}=1-\exp[-\tau_{\lambda}]$ (see, e.g., Castor, 1970, their §6, Rybicki and Hummer 1983, their §3) such that
@@
where $\tau_{\lambda}$ is given by line absorption in Equations 4 and 5 with integral limits consistent with line emission in Equation 10, $\epsilon$ is the filling factor to account for clumpiness because individual clouds do not sample every possible sightline even when they are isotropically distributed, and $f_{\rm esc,\lambda}$ is the fraction of photons which escape radiative trapping at a given wavelength $\lambda$. The escape fraction of photons depends on the fraction $p_{i}=A_{i}/\sum A$ of photons which are produced by a transition $i$ given by the transition rate $A_{i}$ divided by the total radiative depopulation rate $\sum A$, where each $i$ corresponds to a resonant ($p_{res}$) or fluorescent ($p_{fluor}$) channel. Following formalism from Castor (1970, their Equation A12, see also , their Equation 3.19), the probability that a resonant photon can escape from radiative “trapping” through resonant scattering is given by the ratio of emission to absorption $(1-\exp[-\tau_{\rm res,\lambda}])/\tau_{\rm res,\lambda}$. Trapped photons either scatter through emission and re-absorption in the resonant channel or re-emit in the fluorescent channel. After multiple scatterings, the resonant photons are increasingly trapped relative to the fluorescent photons, meaning that the escape fraction is
@@
for fluorescent photons under the infinite scattering limit given by Scarlata and Panagia (2015, their Equations 20a,b).
@@ 2.2 Solutions to Wind Equations of Motion
Velocity profiles for winds can take on a variety of shapes arising from different treatments of the wind equations of motion (e.g., CAK theory Castor et al., 1975a; Lamers and Morton, 1976; Abbott, 1977; Barlow and Cohen, 1977; Castor and Lamers, 1979). The simple $\beta$ law approximation
@@
first appears in early treatment of Wolf-Rayet P Cygni lines for $\beta=0.5$ (Equation 23 Castor, 1970) and arises from a generalization of CAK theory solutions (see, e.g., Equation 47 of Castor et al., 1975a) without loss of physical insight (see, e.g., Barlow and Cohen, 1977; Castor and Lamers, 1979; Pauldrach et al., 1986).The CAK $\beta$ law approximation is ubiquitous in the study of stellar winds (e.g., Vink et al., 2001; Vink and Sander, 2021; Telford et al., 2024) and has even seen application to galactic outflows (e.g., Chisholm et al., 2016; Flury et al., 2023). Given its widespread adoption and underlying physical motivation, the CAK $\beta$ law is our default and fiducial velocity field.
In addition to the $\beta$ law approximation to CAK theory, OutLines includes two alternate solutions to the wind equations proposed in the literature. The first, from Steidel et al. (2010), assumes a power law approximation to the radial acceleration field to obtain
@@
where the normalization terms are implicit in $w$ and $x$ and $r_{min}=1$. While those authors do not provide a corresponding physical model, their acceleration power law has a velocity field qualitatively similar to those of Ipavich (1975), which assumed cosmic ray driving, and Chevalier and Clegg (1985), which assumed constant momentum and energy deposition into a steady-state wind from a fluid in which it is entrained (see also, e.g., Owocki, 2025), and formally akin to radiative acceleration of expanding clouds (see Equations 16-17 Mathews, 1974). Note that the acceleration power law field with $\beta=2$ is identical to that of the CAK $\beta$ law when $\beta=0.5$.
The second velocity field alternative, from various sources (e.g., Faucher-Giguère and Quataert, 2012), constitutes a power law approximation to the radial velocity field such that
@@
where the wind is assumed to accelerate from rest at $R_{0}$ and $A$ scales the field. We set $A=0.5$ to allow clouds to reach 50% terminal velocity in roughly the same range of radii as the CAK theory $\beta$ law and generally overlap with the same region in $w$-$x$ space as both the CAK $\beta$ law and acceleration power law fields. Single-shell models of stellar wind bubbles can give rise to power law velocity fields from the shared time dependence of $r\propto t^{3/5}$ and $v\propto t^{-2/5}$ ($v\propto r^{-2/3}$, e.g., Weaver et al., 1977; Dyson, 1977). Building on this framework, Faucher-Giguère and Quataert (2012) arrive at a velocity power law solution assuming continuous production of “shells” by an AGN to produce a wind.
$w_{\rm nfr}=w[x|\beta]\left(1-\frac{v_{0}}{v_{\infty}}\right)+\frac{v_{0}}{v_{\infty}}$ (18)
(e.g., Castor and Lamers, 1979). Emergent profiles are similar to those presented in Tenorio-Tagle et al. (1996), which included distributions of non-zero shell ages, and to those presented in Scarlata and Panagia (2015), which account for non-zero launch velocities. For visualization purposes, we assume in example profiles in §2 that $v_{0}=0.1v_{\infty}$.
@@
Figure 1 depicts the different velocity fields for a range of $\beta$. From this comparison, as suggested by the formalism, the power law velocity field approaches $v_{\infty}$ more rapidly than the power law acceleration field or the $\beta$-law approximation to CAK theory, reaching $v_{\infty}$ at distances far closer to the launch radius. Moreover, the power law velocity field does not approach $v_{\infty}$ asymptotically as do the other velocity fields, meaning that the stopping condition at $v_{\infty}$ is imposed solely by integral limits rather than the formalism for the velocity field. We note that physical wind models almost always exhibit an acceleration which gradually decreases as the outflow asymptotically approaches some terminal velocity (e.g., Castor et al., 1975a; Barlow and Cohen, 1977; Chevalier and Clegg, 1985; Thompson et al., 2015; Krumholz et al., 2017; Ishibashi et al., 2018; Quataert et al., 2022b), indicating that the underlying wind physics strongly favors a CAK $\beta$ law or acceleration power law field over a velocity power law field. While individual wind bubble models indicate a velocity power law field (e.g., Weaver et al., 1977; Dyson, 1977), in ensembles of shells, the distribution of shell ages will dominate the net velocity field (e.g., Tenorio-Tagle et al., 1996) and is likely determined by the physical driving mechanism (e.g., Chu and Kennicutt, 1994), suggesting that the CAK theory or acceleration power law velocity fields provide a more physically appropriate description for wind scenarios. That being said, the velocity power law field prescription can be consistent with turbulent streaming including the classic Kolmogorov cascade, with the case of $\beta=1$ indicative of Couette flow (e.g., Binette et al., 2009), although such scenarios may not capture the underlying driving mechanism.
@@
Figure 1: Radial velocity fields included in OutLines and given by Equations 15-17. Acceleration from rest at the base $R_{0}$ of the wind ($x=1$) proceeds until the wind reachs the terminal velocity $v_{\infty}$ ($w[x]=1$). Velocity fields include the Pauldrach et al. (1986) $\beta$ law approximation to CAK theory (Castor et al., 1975a) (blue, Equation 15), the Steidel et al. (2010) power law acceleration (green, Equation 16), and the power law velocity field (purple, Equation 17). Dash length increases with increasing $\beta$.